# Generative modeling of density regression through tree flows

### v1

@@ 3.1 Real-world tasks with univariate outcomes
We compare the performance of our proposed method with other methods for conditional density estimation on these univariate tasks. Specifically, we assess the log likelihood of the test set using our method against various established methods including NGBoost (Duan et al., 2020), PGBM (Sprangers et al., 2021), RoNGBa (Ren et al., 2019), TreeFlow (Wielopolski and Zięba, 2023), enhanced versions of KMN and MDN (“KMN+” and “MDN+”) (Rothfuss et al., 2019), Bayesian radial normalizing flows (RNF) (Trippe and Turner, 2018), Bayesian neural networks with homoscedastic Gaussian likelihoods using a mean-field variational approximation (MF), Mixture Density Networks (MDN) (Papamakarios and Murray, 2016), neural networks with latent variable inputs (LV) (Depeweg et al., 2017), Bayesian neural networks with homoscedastic Gaussian likelihoods using Hamiltonian Monte Carlo (HMC) Bui et al. (2016), and Bayesian neural networks with dropout (Dropout) (Gal and Ghahramani, 2016). For our method, the algorithm initially fits tree-CDFs with node-level Logistic Regression and a maximum tree depth of 6. It then switches to node-level MLP with hidden layers sized (4,4) with the maximum tree depth reduced to 4. The hyperparameter that controls the level of $l_{1}$ penalty on the imbalanced splits in our method, $\eta$ (detailed in Appendix B), is set to 0.1. For scale-specific shrinkage, we set $c_{0}=0.05$ and $\gamma=0.5$. The detailed experimental specifications and source of results are available in the Appendix F.
@@ 3.4 Data generation
We demonstrate the data generation capabilities of our proposed model using microbiome compositional data from 16S sequencing experiments. The DIABIMMUNE dataset (Kostic et al., 2015) includes microbiome compositions from 777 stool samples collected from 33 infants over a period of three years. For each observation $(x_{i},y_{i})$, the covariate $x_{i}$ is the age at collection, and the outcome $y_{i}$ is the 17-dimensional microbiome composition aggregated at the class level. The outcomes are normalized to relative abundances, i.e., the elements of each $y_{i}$ sum to 1. The proposed model was trained on the full dataset with $c_{0}=0.05$, $\gamma=0.5$, $\eta=0.1$, and maximum depth of the trees is set to 6. No rotations were applied. With the trained model, one sample is generated for each $x_{i}$, mimicking the conditions under which the original data were collected.
Figure 4 displays a principal coordinate analysis (PCoA) of the Bray-Curtis similarity of training and simulated samples. The simulated samples show similar marginal and conditional distributions to the training data, particularly in the lower-dimensional subspaces defined by the first four main axes of the PCoA. For additional visual comparison, a scatterplot matrix of the training and simulated data can be found in Figure 9 in Appendix G.
Figure 4: Principal coordinate analysis (PCoA) of Bray-Curtis similarity of training (upper row) and simulated (lower row) samples. The first 4 PCoA axes account for $85\%$ of the variability in the data. The color of the points indicates the age (in days) of the infant, which is the covariate in this example.
@@ G.1 Effect of flexible splitting and combination of classifiers
G.2 Results shown in Figure 1
@@
boston concrete energy power wine yacht protein kin8nm naval
Ours -2.44$\pm$0.04 -2.72$\pm$0.04 -0.72$\pm$0.03 -2.66$\pm$0.01 1.98$\pm$0.06 -0.45$\pm$0.07 -2.20$\pm$0.01 1.05$\pm$0.01 6.29$\pm$0.01
NGBoost -2.43$\pm$0.15 -3.04$\pm$0.17 -0.60$\pm$0.45 -2.79$\pm$0.11 -0.91$\pm$0.06 -0.20$\pm$0.26 -2.81$\pm$0.03 0.49$\pm$0.02 5.34$\pm$0.04
PGBM -2.67$\pm$0.10 -2.75$\pm$0.21 -1.74$\pm$0.04 -2.60$\pm$0.02 -0.97$\pm$0.20 -0.05$\pm$0.28 -2.79$\pm$0.01 0.54$\pm$0.04 3.44$\pm$0.04
RoNGBa -2.48$\pm$0.16 -2.94$\pm$0.18 -0.37$\pm$0.28 -2.65$\pm$0.08 -0.91$\pm$0.08 -1.03$\pm$0.44 -2.76$\pm$0.03 0.60$\pm$0.03 5.49$\pm$0.04
KMN+ -2.38$\pm$0.03 -3.33$\pm$0.01 -1.56$\pm$0.02 -2.88$\pm$0.01 0.61$\pm$0.02 -2.02$\pm$0.03 -2.44$\pm$0.01 0.95$\pm$0.01 3.16$\pm$0.01
MDN+ -2.34$\pm$0.05 -3.15$\pm$0.02 -1.34$\pm$0.01 -2.80$\pm$0.01 0.52$\pm$0.03 -1.84$\pm$0.02 -2.43$\pm$0.01 1.16$\pm$0.01 3.21$\pm$0.01
TreeFlow NA -3.02$\pm$0.15 -0.85$\pm$0.35 -2.65$\pm$0.06 0.56$\pm$0.62 -0.72$\pm$0.40 -2.02$\pm$0.02 1.03$\pm$0.06 5.54$\pm$0.16
Dropout -2.46$\pm$0.25 -3.04$\pm$0.09 -1.99$\pm$0.09 -2.89$\pm$0.01 -0.93$\pm$0.06 -1.55$\pm$0.12 -2.89$\pm$0.01 0.95$\pm$0.01 3.80$\pm$0.01
HMC -2.27$\pm$0.03 -2.72$\pm$0.02 -0.93$\pm$0.01 -2.70$\pm$0.01 -0.91$\pm$0.02 -1.62$\pm$0.02 -2.77$\pm$0.01 1.35$\pm$0.01 7.31$\pm$0.01
LV-15 -2.64$\pm$0.05 -3.06$\pm$0.03 -0.74$\pm$0.03 -2.81$\pm$0.01 -0.98$\pm$0.02 -1.01$\pm$0.04 NA NA NA
LV-5 -2.56$\pm$0.05 -3.08$\pm$0.02 -0.79$\pm$0.02 -2.82$\pm$0.01 -0.96$\pm$0.01 -1.15$\pm$0.05 NA NA NA
MDN-2 -2.65$\pm$0.03 -3.23$\pm$0.03 -1.60$\pm$0.04 -2.73$\pm$0.01 -0.91$\pm$0.04 -2.70$\pm$0.05 NA NA NA
MDN-20 -2.74$\pm$0.03 -3.27$\pm$0.02 -1.48$\pm$0.04 -2.68$\pm$0.01 1.21$\pm$0.06 -2.76$\pm$0.07 NA NA NA
MDN-5 -2.73$\pm$0.04 -3.28$\pm$0.03 -1.63$\pm$0.06 -2.70$\pm$0.01 1.43$\pm$0.07 -2.54$\pm$0.10 NA NA NA
MF -2.62$\pm$0.06 -3.00$\pm$0.03 -0.57$\pm$0.04 -2.79$\pm$0.01 -0.97$\pm$0.01 -1.00$\pm$0.10 NA NA NA
RNF-2 -2.40$\pm$0.06 -3.03$\pm$0.05 -0.44$\pm$0.04 -2.73$\pm$0.01 -0.87$\pm$0.02 -0.30$\pm$0.04 NA NA NA
RNF-5 -2.37$\pm$0.04 -2.97$\pm$0.03 -0.67$\pm$0.15 -2.68$\pm$0.01 -0.76$\pm$0.10 -0.21$\pm$0.09 NA NA NA
G.3 Additional results
@@
G.4 Additional figures
This section contains additional figures for the experiments. Specifically, the ground truth and estimated density for the simulation examples are provided in  6-8, and the scatterplot matrix of training data and simulated data in the DIABIMMUNE example is visualized in Figure 9 in Appendix G.
@@ G.4 Additional figures
Figure 9: Training data and simulated data for the microbiome data generation in section 3.4.

### v2

@@ 3.1 Real-world tasks with univariate outcomes
We compare the performance of our proposed method with other methods for conditional density estimation on these univariate tasks. Specifically, we assess the log likelihood of the test set using our method against various established methods including NGBoost (Duan et al., 2020), PGBM (Sprangers et al., 2021), RoNGBa (Ren et al., 2019), TreeFlow (Wielopolski and Zięba, 2023), enhanced versions of KMN and MDN (“KMN+” and “MDN+”) (Rothfuss et al., 2019), Bayesian radial normalizing flows (RNF) (Trippe and Turner, 2018), Bayesian neural networks with homoscedastic Gaussian likelihoods using a mean-field variational approximation (MF), Mixture Density Networks (MDN) (Papamakarios and Murray, 2016), neural networks with latent variable inputs (LV) (Depeweg et al., 2017), Bayesian neural networks with homoscedastic Gaussian likelihoods using Hamiltonian Monte Carlo (HMC) Bui et al. (2016), and Bayesian neural networks with dropout (Dropout) (Gal and Ghahramani, 2016). For our method, the algorithm initially fits tree-CDFs with node-level Logistic Regression and a maximum tree depth of 6. It then switches to node-level MLP with hidden layers sized (4,4) with the maximum tree depth reduced to 4. The hyperparameter that controls the level of $l_{1}$ penalty on the imbalanced splits in our method, $\eta$ (detailed in Appendix B), is set to 0.1. For scale-specific shrinkage, we set $c_{0}=0.05$ and $\gamma=0.5$. The detailed experimental specifications and source of results are available in the Appendix F. As an ablation experiment of the scale-specific shrinkage rate, we set $\gamma=0$ and held all other hyperparameter values the same as the current setting, and The detailed experimental specifications and source of results are available in the Appendix F.
@@
To quantify the impact of scale-specific learning rates, we conducted an ablation experiment by setting $\gamma=0$ (holding all other hyperparameters unchanged). As shown in table 6 in Appendix G, the proposed method with scale-specific learning rates (with $\gamma>0$) outperforms that with a constant learning rate on most of the datasets.
@@ 3.4 Data generation
We demonstrate the data generation capabilities of our proposed model using microbiome compositional data from 16S sequencing experiments. The DIABIMMUNE dataset (Kostic et al., 2015) includes microbiome compositions from 777 stool samples collected from 33 infants over a period of three years. For each observation $(x_{i},y_{i})$, the covariate $x_{i}$ is the age at collection, and the outcome $y_{i}$ is the microbiome composition at the operational taxonomic unit (OTU) level. The outcomes are normalized to relative abundances, i.e., the elements of each $y_{i}$ sum to 1. We keep the 100 OTUs with the highest relative abundance. The proposed model was trained on the full dataset with $c_{0}=0.1$, $\gamma=0.5$, $\eta=0.1$, and maximum depth of the trees is set to 4. No rotations were applied. With the trained model, one sample is generated for each $x_{i}$, mimicking the conditions under which the original data were collected.
Figure 4 displays a principal coordinate analysis (PCoA) of the Bray-Curtis similarity of training and simulated samples. The simulated samples show similar marginal and conditional distributions to the training data, particularly in the lower-dimensional subspaces defined by the first four main axes of the PCoA.
Figure 4: Principal coordinate analysis (PCoA) of Bray-Curtis similarity of training (upper row) and simulated (lower row) samples. The color of the points indicates the age (in days) of the infant, which is the covariate in this example.
@@ G.1 Effect of flexible splitting and combination of classifiers
G.2 Effect of scale-specific shrinkage rates
The proposed method with scale-specific learning rates ($\gamma=0.5$) and fixed learning rates ($\gamma=0$) are compared in Table 6. The other hyperparameters are set to the same values as in Section 3.1.
Table 6: Comparison of predictive scores for different datasets with $\gamma=0$ and $\gamma=0.5$
Dataset $\gamma=0$ $\gamma=0.5$
kin8nm 0.99 $\pm$ 0.01 1.05 $\pm$ 0.01
bostonHousing -2.53 $\pm$ 0.05 -2.44 $\pm$ 0.04
power-plant -2.68 $\pm$ 0.01 -2.66 $\pm$ 0.01
concrete -2.78 $\pm$ 0.05 -2.72 $\pm$ 0.04
protein-tertiary-structure (dequantized) -2.13 $\pm$ 0.01 -2.13 $\pm$ 0.01
yacht -0.56 $\pm$ 0.10 -0.45 $\pm$ 0.07
naval-propulsion-plant 6.48 $\pm$ 0.01 6.29 $\pm$ 0.01
G.3 Results shown in Figure 1
@@
boston concrete energy power yacht protein kin8nm naval
Ours -2.44$\pm$0.04 -2.72$\pm$0.04 -0.72$\pm$0.03 -2.66$\pm$0.01 -0.45$\pm$0.07 -2.20$\pm$0.01 1.05$\pm$0.01 6.29$\pm$0.01
NGBoost -2.43$\pm$0.15 -3.04$\pm$0.17 -0.60$\pm$0.45 -2.79$\pm$0.11 -0.20$\pm$0.26 -2.81$\pm$0.03 0.49$\pm$0.02 5.34$\pm$0.04
PGBM -2.67$\pm$0.10 -2.75$\pm$0.21 -1.74$\pm$0.04 -2.60$\pm$0.02 -0.05$\pm$0.28 -2.79$\pm$0.01 0.54$\pm$0.04 3.44$\pm$0.04
RoNGBa -2.48$\pm$0.16 -2.94$\pm$0.18 -0.37$\pm$0.28 -2.65$\pm$0.08 -1.03$\pm$0.44 -2.76$\pm$0.03 0.60$\pm$0.03 5.49$\pm$0.04
KMN+ -2.38$\pm$0.03 -3.33$\pm$0.01 -1.56$\pm$0.02 -2.88$\pm$0.01 -2.02$\pm$0.03 -2.44$\pm$0.01 0.95$\pm$0.01 3.16$\pm$0.01
MDN+ -2.34$\pm$0.05 -3.15$\pm$0.02 -1.34$\pm$0.01 -2.80$\pm$0.01 -1.84$\pm$0.02 -2.43$\pm$0.01 1.16$\pm$0.01 3.21$\pm$0.01
TreeFlow NA -3.02$\pm$0.15 -0.85$\pm$0.35 -2.65$\pm$0.06 -0.72$\pm$0.40 -2.02$\pm$0.02 1.03$\pm$0.06 5.54$\pm$0.16
Dropout -2.46$\pm$0.25 -3.04$\pm$0.09 -1.99$\pm$0.09 -2.89$\pm$0.01 -1.55$\pm$0.12 -2.89$\pm$0.01 0.95$\pm$0.01 3.80$\pm$0.01
HMC -2.27$\pm$0.03 -2.72$\pm$0.02 -0.93$\pm$0.01 -2.70$\pm$0.01 -1.62$\pm$0.02 -2.77$\pm$0.01 1.35$\pm$0.01 7.31$\pm$0.01
LV-15 -2.64$\pm$0.05 -3.06$\pm$0.03 -0.74$\pm$0.03 -2.81$\pm$0.01 -1.01$\pm$0.04 NA NA NA
LV-5 -2.56$\pm$0.05 -3.08$\pm$0.02 -0.79$\pm$0.02 -2.82$\pm$0.01 -1.15$\pm$0.05 NA NA NA
MDN-2 -2.65$\pm$0.03 -3.23$\pm$0.03 -1.60$\pm$0.04 -2.73$\pm$0.01 -2.70$\pm$0.05 NA NA NA
MDN-20 -2.74$\pm$0.03 -3.27$\pm$0.02 -1.48$\pm$0.04 -2.68$\pm$0.01 -2.76$\pm$0.07 NA NA NA
MDN-5 -2.73$\pm$0.04 -3.28$\pm$0.03 -1.63$\pm$0.06 -2.70$\pm$0.01 -2.54$\pm$0.10 NA NA NA
MF -2.62$\pm$0.06 -3.00$\pm$0.03 -0.57$\pm$0.04 -2.79$\pm$0.01 -1.00$\pm$0.10 NA NA NA
RNF-2 -2.40$\pm$0.06 -3.03$\pm$0.05 -0.44$\pm$0.04 -2.73$\pm$0.01 -0.30$\pm$0.04 NA NA NA
RNF-5 -2.37$\pm$0.04 -2.97$\pm$0.03 -0.67$\pm$0.15 -2.68$\pm$0.01 -0.21$\pm$0.09 NA NA NA
G.4 Additional results
@@
G.5 Additional figures
This section contains additional figures for the experiments. Specifically, the ground truth and estimated density for the simulation examples are provided in  6-8.